OpenStudy (anonymous):

convert the radical below into a fractional exponent /4^3

OpenStudy (anonymous):



OpenStudy (turingtest):

the number of the root goes in the denominator, and the exponent on the base is the numerator:\[\sqrt[b]{x^a}=x^{a/ b}\] here a=3 b=2 (because it is a SQUARE root) x=4 so what's the final answer?